Urea (CH4N2O) is a common fertilizer that can be synthesized by the reaction of ammonia (NH3) with carbon dioxide as follows: 2NH3(aq)+CO2(aq)→CH4N2O(aq)+H2O(l) In an industrial synthesis of urea, a chemist combines 138.7 kg of ammonia with 211.4 kg of carbon dioxide and obtains 167.4 kg of urea.


Determine the theoretical yield of urea

Express your answer using four significant figures.


1
Expert's answer
2021-05-31T02:24:47-0400

Theoretical yield of Urea can be calculated as

"\frac{(138.7 \; kg \; NH_3) }{ (17.04 \; g \;NH_3/mol)} = 8.1396 \;kmol \; NH_3 \\

\frac{(211.4 \;kg \;CO_2) }{ (44.01\; g \;CO_2/mol)} = 4.80345 \;kmol \;CO_2"

8.1396 kilomoles of NH3 would react completely with 8.1396 "\times \frac{1}{2}" = 4.06983 kilomoles of CO2, but there is more CO2 present than that, so CO2 is in excess and NH3 is the limiting reactant.

(8.1396 kmol NH3) "\times" (1 mol CH4N2O / 2 mol NH3) "\times" (60.07 g/mol CH4N2O) = 244.47 kg CH4N2O in theory

Answer: Theoretical yield of Urea is 244.47 kg